A piano wire 0.5m long and mass 5gm is streched by a tension of 400N . What is the number of highest overtone that can be heared by a person?

Grade:12

1 Answers

Vikas TU
7 years ago
Apply Formulae:
F = (2k+1)v/4l
v = root(T/n)
where n is mass per unit length .i.e. mass density.
n = 0.005/0.5 = 0.01
v =  root(400/0.01) = 200m/s
F = (2k + 1)200/4*0.5 = > (2k+1)100